### 10 Apr 2026 v2

# GIST: Gauge-Invariant Spectral Transformers for Scalable Graph Neural Operators

Abstract
Adapting transformer positional encodings to graphs and meshes faces a fundamental tension: exact spectral methods require cubic-complexity eigendecomposition and inadvertently break gauge invariance through numerical solver artifacts, while existing efficient approximations sacrifice gauge symmetry by design. Both failure modes cause catastrophic generalization loss in inductive settings, where models fail when encountering different spectral decompositions of similar graphs or different discretizations of the same domain. We propose GIST (Gauge-Invariant Spectral Transformer), a graph transformer resolving this tension by restricting attention to pairwise inner products of efficient approximate spectral embeddings. We prove these inner products estimate an exactly gauge-invariant graph kernel at end-to-end $\mathcal{O}(N)$ complexity, and establish a formal connection between gauge invariance and discretization invariance: gauge invariance guarantees discretization-invariant learning with bounded mismatch error, making GIST the first scalable graph transformer with provable neural operator guarantees. Empirically, GIST matches state-of-the-art on standard graph benchmarks (e.g., achieving 99.50% micro-F1 on PPI) while uniquely scaling to mesh-based neural operator benchmarks with up to 750K nodes, achieving state-of-the-art on the AirfRANS, ShapeNet-Car, DrivAerNet, and DrivAerNet++ benchmarks.
Graph Neural Networks, Transformers, Spectral Methods, Gauge Invariance, Neural Operators
1 Introduction
Following their incredible success for processing sequential data in Natural Language Processing, Transformers (Vaswani2017) have been demonstrating a remarkable capacity for handling data of increasing structural complexity. Lee2019 have proposed a variant of the transformer block for permutation invariant data with their Set Transformer architecture; Dosovitskiy2021 have adapted the self-attention mechanism to 2D images with the very influential Vision Transformer architecture; and Bertasius2021 have extended transformers to video analysis with their Video Vision Transformer, demonstrating how attention mechanisms capture both spatial and temporal dependencies across video frames. This progression from sequential text to increasingly structured data suggests that Transformers are poised to tackle even more complex data structures, including irregular meshes and graphs.
Indeed, recent developments in adapting Transformers to graphs have shown promising results in capturing global relationships through self-attention, overcoming the long-range dependency limitations of GNNs that rely on localized message passing (Dwivedi2022; Zhu2023c).
Two Barriers to Scalable Graph Transformers.
However, adapting transformers to graphs introduces two distinct barriers that have not been simultaneously addressed. First, there is a computational barrier: exact spectral graph embeddings, while theoretically natural, require eigendecomposition of the graph Laplacian, scaling as $\mathcal{O}(N^{3})$ for dense graphs or $\mathcal{O}(N^{2})$ for sparse graphs (where $N$ is the number of nodes), which is prohibitively expensive for large-scale graphs. Second, there is a gauge invariance barrier: both exact and approximate spectral methods may inadvertently break gauge invariance, i.e. the inherent freedom to rotate eigenvectors, flip signs, or choose among degenerate eigenvectors. Exact methods may break it through numerical solver artifacts (sign choices, eigenspace ordering, degeneracy handling, see Bronstein2017), while approximate methods commit to a specific basis decomposition through their approximation scheme, sacrificing freedom in spectral choices. This gauge invariance breaking introduces spurious inductive biases tied to arbitrary numerical choices, causing models trained with one set of choices to fail catastrophically when evaluated with different spectral decompositions or numerical solvers, particularly in inductive learning tasks where models must generalize to unseen graph structures.
The Gauge Invariance Challenge.
Spectral embeddings depend on arbitrary choices: eigenvector signs, basis orderings, and solver artifacts. A neural network trained on embeddings from one set of choices will learn features that do not transfer when applied with different choices or to different graphs, fundamentally breaking generalization in inductive learning. This gauge dependence is particularly problematic when moving between different graph representations or numerical solvers.
This problem becomes critical for neural operators, where the input domain (a physical system or geometric shape) is discretized into a computational mesh. Different mesh resolutions induce different graph Laplacians and spectral decompositions, each with distinct gauge choices. Models trained on one discretization cannot transfer to another because spectral embeddings at different resolutions are incomparable and the attention kernels cannot be meaningfully compared across discretizations. Gauge invariance, making representations and learned operators independent of arbitrary spectral choices, is essential for both robust inductive learning and for constructing neural operators with provably bounded discretization error.
Existing approaches address only one barrier at a time: (1) spectral methods like SAN (Kreuzer2021) maintain gauge invariance but require full eigendecomposition; (2) approximate spectral methods achieve linear complexity but sacrifice gauge invariance; (3) generic linear transformers reduce attention complexity but ignore graph structure. Recent spectral invariant methods achieve gauge invariance through careful attention design, but still require full eigendecomposition ($\mathcal{O}(N^{3})$ complexity) or quadratic attention ($\mathcal{O}(N^{2})$), limiting practical scalability. Moreover, they lack discretization error controls needed for neural operator applications where consistency across mesh resolutions is critical.
Our Contribution: GIST.
We propose Gauge-Invariant Spectral Transformers (GIST), which overcomes both barriers through a key insight: gauge invariance can be preserved by restricting attention to inner products between spectral embeddings, which remain invariant even under approximate spectral computations. We show that a broad class of efficient $\mathcal{O}(N)$ spectral embeddings of the form $\widetilde{\Phi}=f(P)\cdot R$, where $f$ is any spectral filter and $R\in\mathbb{R}^{N\times r}$ is a random matrix with $\mathbb{E}[RR^{\top}]=\mathds{1}$, share a key property: their inner products are unbiased estimators of the graph kernel $K=f(P)f(P)^{\top}$, which depends only on the graph operator $P$ and is therefore exactly gauge-invariant. By restricting attention operations to these inner products and combining with linear attention, we achieve end-to-end $\mathcal{O}(N)$ complexity while recovering gauge invariance exactly at the population level, with statistical error $\mathcal{O}(r^{-1/2})$. Unlike prior approaches that address these barriers separately, GIST achieves gauge invariance and linear complexity simultaneously. Our contributions in short are:
  1. 1.
    identifying gauge invariance breaking as a failure mode shared by both exact and approximate spectral methods, and demonstrating that it causes catastrophic generalization loss in inductive graph settings;
  2. 2.
    proving that gauge invariance is necessary for discretization invariance with bounded discretization error and establishing the first discretization mismatch bound for a graph neural operator;
  3. 3.
    proposing GIST, the first graph transformer that simultaneously achieves gauge invariance, end-to-end $\mathcal{O}(N)$ complexity, and global spectral expressivity, enabling scalable neural operators with provable error bounds validated by state-of-the-art results on mesh problems with up to 750K nodes.
2 Related Works
Graph Transformers.
Graphormer (Ying2021) introduces the idea of integrating structural encodings such as shortest path distances and centrality in Transformers. Similarly, Dwivedi2022 propose LSPE (Learnable Structural and Positional Encodings), an architecture that decouples structural and positional representations. Kreuzer2021 propose Spectral Attention Network (SAN), which introduces learned positional encodings from the full Laplacian spectrum. Park2022a develop Graph Relative Positional Encoding (GRPE), which extends relative positional encoding to graphs by considering features representing node-topology and node-edge interactions. Hierarchical Graph Transformer (Zhu2023c) addresses scalability to million-node graphs through graph hierarchies and coarsening techniques. SpecFormer (bo2023specformer) and PolyFormer (Chen2025) are recent spectral Transformers that improve over SAN by leveraging approximate spectral bases or low-rank polynomial Laplacian filters to enhance scalability and accuracy on graph tasks.
Several recent architectures aim to bridge local and global reasoning via structural encodings and scalable attention. GraphGPS (rampasek2022GPS) adopts a hybrid paradigm that decouples local message passing from global attention while integrating spectral positional encodings to capture multi-scale interactions. However, unlike GIST, it lacks a mechanism to algorithmically preserve gauge invariance, leaving the model susceptible to the arbitrary numerical artifacts and sign flips inherent in spectral decompositions. While models like Exphormer (shirzad2023exphormer) attempt to address the resulting complexity by replacing full attention with sparse expander-based mechanisms, they still struggle with the high overhead of global coordinate systems. Similarly, tokenized approaches such as TokenGT (hamilton2017inductive) and NAGphormer (chen2023nagphormer) treat graphs as sets of tokens with [CLS]-style readouts, yet these models scale poorly due to memory-intensive tokenization and positional encoding costs that grow super-linearly with graph size. Consequently, these frameworks are rarely evaluated on large-scale inductive tasks like ogbn-arxiv, where maintaining structural encodings becomes computationally prohibitive.
Scalable Attention Architectures.
Recent advances tackled the quadratic scaling of self-attention through various approaches, including cross-attention bottlenecks that map inputs to fixed-size latent representations or concepts (Jaegle2021b; Rigotti2022a), kernel-based attention mechanisms using random feature approximations (Choromanski2020), feature map decomposition methods that linearize the attention computation (Katharopoulos2020), and memory-efficient variants with sub-linear complexity (Likhosherstov2021). As noted by Dao2024, many such linear transformer models are directly related to linear recurrent models such as state-space-models (Gu2021; Gu2022b; Gu2023a; ChennuruVankadara2024)
Neural Operators.
Further addressing the scalability of these graph-based methods is essential for applying them to complex domains such as geometry meshes and point clouds. In these settings, graphs are induced by the connectivity of an underlying continuous object whose discretization is not unique: it can be sampled at arbitrarily many densities and resolutions. High-density discretizations can render the graph prohibitively large, undermining both efficiency and scalability in existing methods. As a result, efficient mesh downsampling and/or re-discretization onto regular lattices (e.g., via SDF-based volumetric grids), and task-aware coarsening learned by GNNs, were commonly required to make these problems tractable.
Neural operators learn maps between continuous function spaces rather than fixed-dimensional vectors. Two properties are crucial here: (i) discretization invariance, i.e., a single set of parameters applies across discretizations (meshes, resolutions, and sampling locations) of the same underlying continuum problem; and (ii) global integration, i.e., the ability to represent nonlocal interactions via learned integral kernels, rather than being limited to finite-receptive-fields. Formally, a neural operator composes learned integral operators with pointwise nonlinearities, yielding universal approximation results for continuous nonlinear operators and implementations that share weights across resolutions. Our approach preserves these neural operator properties and improves scalability, allowing it to be applied to these cases (kovachki2023neuraloperator).
Foundational operator families. The Fourier Neural Operator (FNO) parameterizes kernels in the spectral domain and evaluates them with FFT-based spectral convolutions, sharing weights across resolutions and enabling efficient nonlocal interactions on grids (li2020fno). The Graph Neural Operator (GNO) realizes the kernel via message passing, supporting irregular meshes and geometry variation while keeping the learned map discretization-agnostic (li2020gno). Convolutional Neural Operators (CNOs) define continuous convolutions with learnable kernels and interpolation, specifying the operator in the continuum and discretizing only at runtime (raonic2023cno).
Hybrid designs pair geometry-aware encoders with operator layers to handle complex shapes. GINO couples a graph encoder/decoder with a latent FNO on a proxy grid from SDF or point-cloud inputs and shows convergence across large 3D, multi-geometry problems (li2023gino). Encoder–decoder operator learners, such as DeepONet, use a branch network for inputs and a trunk network for coordinate queries, directly supporting heterogeneous sampling (lu2021deeponet); U-NO adds a multi-resolution U-shaped backbone for multiscale effects (rahman2022uno).
Transformers as neural operators. Self-attention is a learned, data-dependent kernel integral, and with suitable positional features can approximate continuous maps on variable-length sets for discretization-invariant operator learning; cross-attention evaluates outputs at arbitrary coordinates (tsai2019kernelattention; yun2019transformersuniversal; lee2019settransformer; jaegle2021perceiverio). Transolver casts PDE operator learning as attention from query coordinates to context tokens built from input fields, yielding resolution-agnostic inference and strong generalization across meshes (wu2024transolver). Recent operator transformers, e.g., GNOT, add geometric normalization and gating to stabilize training on irregular meshes and multi-condition PDEs (hao2023gnot).
Positioning GIST.
Existing graph transformers and scalable attention methods address complementary but not simultaneous challenges. Spectral methods like SAN (Kreuzer2021) leverage the full Laplacian spectrum to maintain theoretical expressiveness but incur significant computational costs from full spectral methods. Approximate spectral methods achieve better scalability but completely forsake gauge invariance, exacerbating generalization failures when arbitrary gauge choice differs between training and testing. Generic linear transformers reduce attention complexity but typically ignore graph structure entirely. GIST uniquely combines graph awareness through spectral embeddings, computational efficiency through random projections, gauge invariance through a modified attention mechanism, and linear attention for end-to-end linear scaling. It also preserves the discretization-invariance and global integration properties needed for neural operator applications on mesh regression, unifying graph learning and continuous function approximation in a single framework.
3 Approach
3.1 Preliminaries
Self-attention and positional encoding.
Given query, key and value representations $q_{i}$, $k_{i}$, $v_{i}$ of $N$ tokens with $i=1,\ldots,N$, self-attention (Vaswani2017) famously computes outputs as a weighted sum of values with attention weights determined by query-key similarities:
$o_{i}=\sum_{j=1}^{N}\alpha_{ij}v_{j},\quad\mbox{where }\alpha_{ij}=\mathrm{softmax}_{j}\left(\frac{q^{\top}_{i}k_{j}}{\sqrt{d}}\right).$ (1)
A key insight (Shaw2018) is that positional information can be injected through relative positional biases: $e_{ij}=\frac{q^{\top}_{i}k_{j}}{\sqrt{d}}+b_{ij}$, where $b_{ij}$ reflects distances between positions. For graphs, this can be generalized by replacing $b_{ij}$ with distance measures reflecting the graph structure.
Graph Laplacian and spectral embeddings.
The (normalized) graph Laplacian $\mathcal{L}=\mathds{1}-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$}} induces a natural metric via the resistance distance: $\Omega(i,j)=(e_{i}-e_{j})^{\top}\mathcal{L}^{\dagger}(e_{i}-e_{j})$, where $e_{i}$ is the $i$th standard basis vector and $\mathcal{L}^{\dagger}$ is the Moore-Penrose pseudoinverse (Klein1993). The resistance distance can be expressed via Laplacian eigenmaps, which satisfy:
[…]
with $\lambda_{k},u_{k}$ being the eigenvalues and eigenvectors of $\mathcal{L}$. These eigenmaps are natural graph positional encodings (Dwivedi2021), but exact computation requires $\mathcal{O}(N^{3})$ eigendecomposition, prohibitive for large graphs.
Approximate spectral embeddings and the gauge invariance problem.
Note from equation 2 that the graph metric is determined by pairwise inner products of eigenmaps, since $\Omega(i,j)=\langle\phi_{i},\phi_{i}\rangle-2\langle\phi_{i},\phi_{j}\rangle+\langle\phi_{j},\phi_{j}\rangle$. This means we do not need the eigenvectors themselves, only an efficient way to estimate this kernel. A broad class of randomized methods achieves this, computing embeddings as
$\widetilde{\Phi}=f(P)\cdot R\;\in\mathbb{R}^{N\times r},$ (3)
where $P$ is a graph operator (e.g., the random-walk matrix $D^{-1}A$ or the normalized Laplacian) with eigendecomposition $P=U\Lambda U^{\top}$, $f$ is a spectral filter ($f(P)=U\,\mathrm{diag}(f(\lambda_{1}),\ldots,f(\lambda_{N}))\,U^{\top}$), $R\in\mathbb{R}^{N\times r}$ is a random matrix with $\mathbb{E}[RR^{\top}]=\mathds{1}$, and $\tilde{\phi}_{i}\in\mathbb{R}^{r}$ is the $i$-th row of $\widetilde{\Phi}$. Laplacian eigenmaps equation 2 are the special case $f(\lambda)=\lambda^{-1/2}$, $P=\mathcal{L}$; polynomial filters, rational filters, and random-walk diffusions also fit this template, with $f(P)\cdot R$ computable via repeated sparse matrix-vector products without forming $f(P)$ explicitly.
A key property shared by all methods in this class is that their pairwise inner products are unbiased estimators of the graph kernel $K(i,j):=[f(P)f(P)^{\top}]_{ij}$:
$\mathbb{E}\!\left[\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle\right]=K(i,j)=[f(P)f(P)^{\top}]_{ij},$ (4)
where the expectation is over the draw of $R$ and follows from $\mathbb{E}[RR^{\top}]=\mathds{1}$; when $P$ is symmetric this reduces to $K(i,j)=\sum_{l}f(\lambda_{l})^{2}\,(u_{l})_{i}(u_{l})_{j}$. Crucially, $K(i,j)$ is a function of $P$ alone and therefore exactly gauge-invariant: it is unchanged by sign flips, eigenspace rotations, and numerical solver choices (Proposition A.5). The estimation error is statistical and vanishes as $\mathcal{O}(1/\sqrt{r})$ (Lemma A.6). In our implementation we use FastRP (Chen2019e), a polynomial filter $f(P)=\sum_{i=1}^{k}P^{i}$ computed via sparse random projections and power iterations at $\mathcal{O}(Nrk)$ cost (see Algorithm 1), but any method in this class works equally well, e.g. graph diffusion kernels (Kondor2002), Chebyshev polynomial filters (Hammond2011), or randomized SVD-based embeddings (Halko2011).
Importantly, exact spectral methods can also introduce gauge dependence through eigensolver artifacts (sign choices, eigenspace ordering), making gauge invariance a concern across both exact and approximate approaches (Bronstein2017).
Motivation for gauge-invariant operations.
While the inner products of approximate spectral embeddings estimate a gauge-invariant kernel, the embeddings themselves depend on arbitrary choices: the random matrix $R$, eigensolver conventions, eigenvalue ordering. Any model operating on these embeddings directly, rather than on their inner products, will learn features that don’t transfer across graphs, resulting in a generalization failure. Our approach addresses this by using approximate spectral embeddings as positional encodings, but restricting the Transformer to operate only on their gauge-invariant inner products.
3.2 Our Approach: GIST.
The key insight is that spectral embeddings of the form $\widetilde{\Phi}=f(P)\cdot R$ produce inner products $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle$ that are unbiased estimators of the exactly gauge-invariant kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ (equation 4). By restricting all operations to depend on the embeddings only through these inner products, gauge invariance is preserved exactly at the population level (with only statistical noise of order $\mathcal{O}(1/\sqrt{r})$) while maintaining computational efficiency.
Figure 1: Gauge-Invariant Spectral Transformer. Left: Gauge-Invariant Spectral Self-Attention operates on graph positional embeddings $\tilde{\phi}$ as queries and keys, and node features $x$ as values. The output of the self-attention operation is then combined with $x$ through a residual connection. Limiting $\tilde{\phi}$ to queries and keys preserves gauge invariance across the self-attention block. The diagram uses standard softmax attention for clarity of exposition; the actual GIST implementation replaces it with linear attention. Right: Gauge-Invariant Spectral Self-Attention is embedded in a Multi-Scale Gauge-Invariant Spectral Transformer Block which comprises 3 parallel branches inspired by EfficientViT.
Gauge-Invariant Spectral Self-Attention.
We now introduce our main contribution: Gauge-Invariant Spectral Transformer (GIST). The first ingredient of GIST is Gauge-Invariant Spectral Self-Attention, which operates on approximate spectral embeddings $\tilde{\phi}_{i}\in\mathbb{R}^{r}$ (the $i$-th row of $\widetilde{\Phi}=f(P)\cdot R$) but restricts attention computations to inner products between embeddings. The key observation is that while the individual embeddings $\tilde{\phi}_{i}$ depend on the random matrix $R$, their inner products $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle$ are unbiased estimators of the gauge-invariant kernel $K(i,j)$ (equation 4), which depends only on $P$. Attention weights computed from these inner products are therefore gauge-invariant and transfer across graphs.
Formally, for each node $i=1,\ldots,N$, Gauge-Invariant Spectral Self-Attention modifies the standard self-attention mechanism as follows:
$q_{i}=\tilde{\phi}_{i},\quad k_{i}=\tilde{\phi}_{i},\quad v_{i}=W_{v}x_{i},$
where $W_{v}$ is a learned linear projection of the node features $x_{i}$, as usual in transformers. This ensures attention logits are based on inner products:
$e_{ij}=q^{\top}_{i}k_{j}=\tilde{\phi}^{\top}_{i}\tilde{\phi}_{j}=K(i,j)+\mathcal{O}(1/\sqrt{r}),$
where $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ is the exactly gauge-invariant kernel from equation 4 (see Fig. 1, left). By limiting the embeddings to the query-key computation and not using them as values, we ensure that downstream layers (which operate on node features) cannot access gauge-dependent information. Algorithm 1 in Section B gives our specific implementation of the spectral embedding computation (which uses FastRP), and Algorithm 2 details the attention implementation. Notice that the softmax in equation 1 is replaced by linear attention in our implementation (see the Linear Self-Attention paragraph below).
Gauge-Equivariant Spectral Self-Attention.
The Gauge-Invariant Spectral Self-Attention operation thus preserves gauge invariance, but at the cost of giving up a lot of the flexibility of regular self-attention. In particular, there is no mechanism that allows for a modification of the vectors $\tilde{\phi}_{i}$ through learning. In fact, applying even just a linear operation on $\tilde{\phi}_{i}$ would break again gauge invariance. However, notice that rescaling each $\tilde{\phi}_{i}$ by a scalar possibly depending on the node features $s(x_{i})\in\mathbb{R}$ would modify the similarity between graph positional embeddings in the same way across gauge choices, since scalars commute with orthogonal projections, meaning that it is an equivariant operation across gauges: $(s(x_{i})\tilde{\phi}_{i})^{\top}(s(x_{j})\tilde{\phi}_{j})=s(x_{i})s(x_{j})(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})=s(x_{i})s(x_{j})\cdot K(i,j)+\mathcal{O}(1/\sqrt{r})$.
Remarkably, such a gauge-equivariant operation can also be straight-forwardly implemented via a modification of self-attention by modifying equation 1 as follows:
$q_{i}=W_{q}x_{i},\quad k_{i}=W_{k}x_{i},\quad v_{i}=\tilde{\phi}_{i},$
and the output in equation 1 such that it is constrained to operate on $\tilde{\phi}_{i}$, i.e. $\tilde{\phi}_{i}^{l+1}=\sum_{j=1}^{N}\alpha_{ij}v_{j}$, where $\tilde{\phi}_{i}^{l+1}$ indicates the graph positional encoding that will be used in the next layer $l+1$. Algorithm 3 in Section B details how the implementation of Gauge-Equivariant Spectral Self-Attention relates to regular Self-Attention.
Linear Self-Attention, and Multi-Scale Architecture.
Gauge-Invariant Spectral Self-Attention ensures that we can compute reliable graph positional encoding with linear time complexity in the number of nodes in the graph $N$. In order to maintain that linear scaling end-to-end, all GIST attention blocks replace softmax attention (introduced above and depicted in Fig. 1, left, for clarity) with the linear transformer by Katharopoulos2020, using a feature map $\varphi$ so that $\alpha_{ij}\propto\langle\varphi(q_{i}),\varphi(k_{j})\rangle$.
Crucially, as feature map we use $\varphi(x)=\mathrm{ReLU}(x)$, which is a map that induces a kernel $k_{0}(\cdot)$ corresponding to the arc-cosine kernel (Cho2009). More specifically, for random features $\tilde{\phi}_{i},\tilde{\phi}_{j}\in\mathbb{R}^{r}$, the attention weights $\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle\approx k_{0}(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})$ converge to a kernel function that depends only on the inner product $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}$. Since $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}$ estimates the exactly gauge-invariant kernel $K(i,j)$ (equation 4), this preserves gauge invariance: attention weights depend only on gauge-invariant quantities. For further considerations on the choice of the feature map $\varphi(\cdot)$ see the note in Appendix B.
To mitigate linear attention’s drawbacks (notably lack of sharp attention scores compared to softmax attention), we design a parallel architecture inspired by EfficientViT (Cai2022), a multi-scale linear attention architecture. Our Multi-Scale Gauge-Invariant Spectral Transformer Block comprises three parallel branches: a feature branch (linear transformer on node features $x$), a local branch (graph convolution on $x$ followed by linear transformer), and a global branch (Gauge-Invariant Spectral Self-Attention followed by Gauge-Equivariant Spectral Self-Attention operating on both $x$ and $\tilde{\phi}$, then linear transformer). The local branch emphasizes local information that linear attention would otherwise diffuse, while the global branch integrates global graph information. See Figure 1 (right panel) for the architecture. Ablation studies in Appendix C.3 confirm that each branch meaningfully contributes to performance.
Complexity Scaling Analysis.
GIST computational complexity is dominated by two components. First, spectral embedding computation scales as $\mathcal{O}(N\cdot r\cdot k)$ where $r$ is the embedding dimension and $k$ is the number of power iterations (e.g., $k$ polynomial filter steps for FastRP). Second, linear transformer blocks with Gauge-Invariant Spectral Self-Attention on $d$-dimensional node features scale as $\mathcal{O}(N\cdot d^{2})$. Overall, this gives end-to-end $\mathcal{O}(N\cdot d^{2}+N\cdot r\cdot k)$ scaling, i.e. linear in the number of nodes $N$. This contrasts with $\mathcal{O}(N^{3})$ for exact eigendecomposition and $\mathcal{O}(N^{2}d)$ for standard quadratic attention. This linear scaling is empirically validated in Appendix C.2: Figure 3 demonstrates that both VRAM consumption and forward pass time grow linearly with graph size up to 500K nodes on DrivAerNet samples, confirming our theoretical analysis and enabling the large-scale neural operator experiments in Section 4.
4 Results
4.1 Theoretical Guarantees: GIST as Discretization-Invariant Neural Operator
We now formalize how gauge invariance enables discretization-invariant learning with bounded error. Physical problems (computational fluid dynamics, structural mechanics, shape analysis) are discretized into computational meshes (i.e. graphs), with different resolutions producing different graph Laplacians and spectral decompositions involving arbitrary gauge choices (sign flips, eigenspace rotations, solver artifacts). Without gauge invariance, parameters fail to transfer across discretizations; gauge invariance ensures attention weights converge to a well-defined continuum kernel with bounded discretization mismatch error.
Proposition (Informal).
Outputs of GIST applied to different discretizations of the same $m$-dimensional manifold, obtained by random sampling nodes, converge to each other with error $\mathcal{O}(n^{-1/(m+4)})+\mathcal{O}(r^{-1/2})$, where $n$ is the coarser resolution and $r$ is the spectral embedding dimension. This ensures learned parameters transfer across arbitrary mesh resolutions with bounded error that vanishes as resolution and embedding dimension increase.
Proof sketch. The result follows from: (1) spectral convergence of graph Laplacians to manifold eigenfunctions at rate $\mathcal{O}(n^{-1/(m+4)})$ (BelkinNiyogi2008; CalderGarciaTrillos2020); (2) spectral embeddings $\widetilde{\Phi}=f(P)\cdot R$ produce inner products that estimate the exactly gauge-invariant kernel $K=f(P)f(P)^{\top}$ with statistical error $\mathcal{O}(r^{-1/2})$ (Lemma A.4, Proposition A.5, Lemma A.6); (3) gauge-invariant attention depends only on these inner products, ensuring parameter transfer with total error $\mathcal{O}(n^{-1/(m+4)})+\mathcal{O}(r^{-1/2})$. Full formulation and proof are found in Appendix A.
This result distinguishes GIST from prior spectral methods: while approximate methods achieve computational efficiency, they lack gauge invariance and thus cannot provide discretization error bounds; while exact methods could maintain invariance, their $\mathcal{O}(N^{3})$ cost prevents scaling to high-resolution meshes where such bounds are most valuable. GIST uniquely combines gauge invariance with end-to-end linear complexity, achieving both theoretical guarantees and practical scalability for neural operator applications.
4.2 Discretization Invariance Empirical Verification
We validate the discretization invariance property predicted by Proposition Proposition with a controlled experiment isolating the effect of gauge-invariance on cross-resolution transfer.
Table 1: Discretization invariance verification. Models trained on 50% decimated mesh and tested on full-resolution mesh of the same car geometry. Gauge-invariant architecture transfers successfully; non-gauge-invariant baseline fails (we show R-squared, averaged over 6 random seeds $\pm$ standard deviation).
Architecture Train $R^{2}$ $\uparrow$ Test $R^{2}$ $\uparrow$
(coarse mesh) (fine mesh)
[…]
Experimental setup.
We train two architectures with matched parameters on a single DrivAerNet (elrefaie2024drivaernet) car at 50% decimation ($\sim$}250K vertices) and test on full resolution ($\sim$}500K vertices) predicting surface pressure. The gauge-invariant architecture is a GIST block followed by a linear transformer block. The non-gauge-invariant baseline uses a linear transformer block with spectral embeddings summed to node features (treating them as standard positional encodings), which breaks gauge invariance by using the projected embeddings directly as features rather than restricting attention to their inner products.
Discretization Invariance Results.
Table 1 shows both architectures achieve essentially the same training performance on the coarse mesh, confirming matched model capacity. However, when evaluated on the fine mesh, the gauge-invariant architecture maintains strong performance (0.840 $R^{2}$, a 15% relative drop), while the non-gauge-invariant baseline fails catastrophically (0.526 $R^{2}$, a 47% relative drop). This controlled experiment directly demonstrates that gauge invariance is essential for parameter transfer across discretizations, validating our theory.
4.3 Node Classification Tasks
To demonstrate the key advantages of GIST, we evaluate it on both transductive and inductive node classification datasets. Transductive tasks are a common graph neural networks paradigm and consist of training and evaluating the model on the same graph, with the goal of predicting at test time node labels that were not provided at training (infilling). Inductive tasks instead operate on a disjoint set of graphs and aim to predict properties on a new graph.
Experiment Setup
We evaluate our method on transductive graph benchmarks using the official training, validation, and test splits and evaluation protocols. For each method, we select optimal hyperparameters by optimizing over the validation split of each dataset. For final results we train on combined training and validation set.
Spectral embedding parameters based on FastRP ($k$ and $r$) are optimized through HPO. Appendix C.1 shows final accuracy is robust to variation, with performance saturating at relatively low $r$, consistent with the $\mathcal{O}(r^{-1/2})$ concentration rate of the kernel estimator (Lemma A.6).
4.3.1 Transductive Tasks
We evaluate our method on the three standard Planetoid citation benchmarks for the transductive setting where the whole graph is observed at train time: Cora (2,708 nodes, 5,429 edges, 1,433 bag-of-words features, seven classes), CiteSeer (3,327 nodes, 4,732 edges, 3,703 features, six classes), and PubMed (19,717 nodes, 44,338 edges, 500 features, three classes). Train-val-test sets follow the Planetoid public split, and we report node-classification accuracy (sen2008collective; yang2016planetoid; kipf2017).
Table 2: Transductive node classification on the Planetoid benchmarks (Cora, Citeseer, Pubmed). We report test accuracy (%) as average $\pm$ standard deviation across random seeds using the standard public split (higher is better). Benchmark results are taken from the following references (a ‘–’ means no reported results): a(Luo2024), b(Choi2022), c(Chen2020) .
Model
Cora
(accuracy $\uparrow$)
Citeseer
(accuracy $\uparrow$)
Pubmed
(accuracy $\uparrow$)
[…]
Although GIST’s gauge invariance is most impactful in inductive and large-scale settings (where spectral gauge changes occur between training and testing), it remains competitive on transductive benchmarks, demonstrating the generality of the architecture. Across these benchmarks, GIST is competitive with strong graph convolutional and transformer-style baselines (see Table 2). On Pubmed, GIST attains the best mean accuracy among the reported methods ($81.20\%\pm 0.41$), narrowly surpassing enhanced GCN variants (e.g., $81.12\%\pm 0.52$) and outperforming GAT/GraphSAGE families. On Cora and Citeseer, GIST achieves results comparable to the top results (within $\sim$1–2 points of GCNII/SGFormer and the enhanced GCN), landing at $84.00\%\pm 0.60$ and $71.31\%\pm 0.50$, respectively.
Table 3: Inductive node classification on PPI, Arxiv, and Photo. Results are reported as micro-F1 (higher is better) as average $\pm$ standard deviation across random seeds. Benchmark results are taken from the following references (a ‘–’ means no reported results): a(Zeng2020), b(Brody2022), c(bo2023specformer), d(Velickovic2018), e(Chen2020), f(Choi2022), g(Chiang2019), h(Chen2025), i(shirzad2023exphormer), j(Wu2024c), k(Ma2024), l(Zhang2018a). source paper reports 95% confidence interval as uncertainty interval which is rescaled to standard deviation. source paper does not specify what the reported uncertainty intervals are.
Model
PPI
(micro-F1 $\uparrow$)
ArXiv
(micro-F1 $\uparrow$)
Photo
(micro-F1 $\uparrow$)
[…]
4.3.2 Inductive Tasks
We evaluate our method on three inductive benchmarks: PPI, Arxiv, and Photo. PPI is a collection of 24 disjoint tissue-specific protein–protein interaction graphs where nodes (proteins) have 50 features and 121 non–mutually-exclusive GO labels. Following the standard split (20 graphs for training, 2 for validation, and 2 for testing), we report micro-averaged F1 on the unseen test graphs. For the Arxiv citation graph (269,343 nodes, 1,166,243 edges, 128 features, and 40 classes) and the Amazon Photo co-purchase network (7,650 nodes, 119,081 edges, 745 features, and 8 classes), we evaluate GIST in the inductive node classification setting using official splits (Fey/Lenssen/2019; Fey/etal/2025).
GIST achieves strong performance across all datasets. On PPI, GIST reaches $99.50\%\pm 0.03$ micro-F1, matching the best large-scale sampling methods and deep residual GCNs (see Table 3, on par with GCNIII and within noise of the strongest GCNII setting ($99.53\%$)). On Arxiv, GIST achieves a mean micro-F1 of $72.12\%\pm 0.21$, competitive with recent spectral transformers (PolyFormer: 72.42%, SpecFormer: 72.37%, Exphormer: 72.44%) while maintaining efficient $\mathcal{O}(N)$ scaling through its linear attention formulation. On the smaller but feature-rich Photo dataset, GIST attains $95.61\pm 0.23$ micro-F1, which slightly surpasses SpecFormer’s state-of-the-art performance of $95.48\pm 0.16$. Note that GraphGPS goes OOM on ArXiv due to poor scalability. These findings demonstrate GIST’s effectiveness as a competitive graph learning approach, validating the successful trade-off between computational overhead and representational power.
4.4 Neural Operators
We validate GIST’s discretization-invariant properties (Section 4.1) on four surface pressure prediction benchmarks of increasing scale, all involving node-level regression of aerodynamic fields on unseen geometries.
AirfRANS (bonnet2022airfrans) is a high-fidelity computational fluid dynamics dataset of 2D airfoils under Reynolds-Averaged Navier-Stokes (RANS) simulation, comprising 1000 airfoil designs with approximately 200K mesh nodes per sample across Reynolds numbers 2–6M and angles of attack from $5^{\circ}$ to $15^{\circ}$. ShapeNet-Car (chang2015shapenet) is a collection of 3D car geometries used for aerodynamic prediction, containing approximately 3700 mesh points per sample. DrivAerNet and DrivAerNet++ (elrefaie2024drivaernet) are distinct high-fidelity CFD datasets of parametric car geometries: DrivAerNet comprises 4,000 designs with approximately 500K surface vertices per car, while DrivAerNet++ extends this to over 10,000 designs with greater geometric diversity across multiple vehicle classes (SUVs, sedans, hatchbacks). All datasets model cars or airfoils as graphs with nodes representing surface mesh vertices and edges following mesh connectivity.
On the intermediate-scale benchmarks (AirfRANS and ShapeNet-Car, Table 4), GIST outperforms Transolver, achieving 0.0028 MSE on AirfRANS (vs. 0.0142) and 4.68 MSE on ShapeNet-Car (vs. 4.99). On the large-scale benchmarks (DrivAerNet and DrivAerNet++, Table 5), GIST achieves state-of-the-art on both: 20.10% relative $\ell_{2}$ error on DrivAerNet (vs. 20.35% for the previous best) and 18.60% on DrivAerNet++ (vs. 20.05%). GIST’s linear scaling enables direct processing of these 500K-node meshes without downsampling, while prior methods require lossy projection to regular grids or lower-dimensional latent spaces. GIST’s spectral attention provides global receptive fields in a single layer, which aids performance on all these tasks.
Table 4: Surface pressure prediction accuracy on AirfRANS and ShapeNet-Car benchmarks. Baseline is Transolver (wu2024transolver). MSE is reported directly. Lower is better.
Model AirfRANS (MSE $\downarrow$) ShapeNet-Car (MSE $\downarrow$)
Transolver 0.0142 4.99
GIST (ours) 0.0028 4.68
Table 5: Surface pressure prediction accuracy on DrivAerNet and DrivAerNet++. Baselines include RegDGCNN (elrefaie2024drivaernet), Transolver (wu2024transolver), FigConvNet (choy2025figconvnet), TripNet (chen2025tripnet), and AdaField (Zouetal2026). MSE is reported in units of $10^{-2}$; Rel L2 is reported as a percentage (%). Lower is better for both metrics.
DrivAerNet DrivAerNet++
[…]
RegDGCNN 9.01 28.49 8.29 27.72
Transolver 5.37 22.52 7.15 23.87
FigConvNet 4.38 20.98 4.99 20.86
TripNet 4.23 20.35 4.55 20.05
GIST (ours) 4.16 20.10 3.63 18.60
5 Conclusions
We presented GIST, a gauge-invariant spectral transformer whose contributions span three levels. First, we identified gauge invariance breaking as a fundamental failure mode shared by both exact and approximate spectral methods, causing catastrophic generalization loss in inductive settings. Second, we proved formally that gauge invariance is a necessary condition for discretization-invariant neural operators, establishing the first discretization mismatch bound $\mathcal{O}(n^{-1/(m+4)})+\mathcal{O}(r^{-1/2})$ for a graph neural operator. Third, we proposed an architecture that achieves gauge invariance, end-to-end $\mathcal{O}(N)$ complexity, and global spectral expressivity simultaneously, three properties that no prior method combines. By restricting attention to inner products of spectral embeddings, which are unbiased estimators of an exactly gauge-invariant kernel, GIST recovers the symmetry algorithmically while maintaining linear scaling.
Empirically, GIST achieves competitive results on standard graph benchmarks (Cora, PubMed, PPI) and sets a new state-of-the-art on large-scale mesh regression (DrivAerNet and DrivAerNet++, up to 500K and 750K nodes, respectively), processing these graphs directly at full resolution where methods requiring quadratic attention or cubic eigendecomposition cannot scale.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning through improved geometric deep learning and neural operator methods. There are many potential societal consequences of our work, and the methods presented could be beneficial for scientific computing applications (e.g., computational fluid dynamics) and could advance the state of graph neural networks.
Appendix A Proofs
A.1 Proof of Proposition Proposition: Gauge-Invariant Spectral Self-Attention
We prove that the Gauge-Invariant Spectral Self-Attention mechanism is discretization-invariant with quantifiable discretization mismatch error through three stages of analysis. The proof establishes that the attention kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ estimated by the spectral embeddings converges to a well-defined continuum kernel as mesh resolution increases, allowing us to bound the discretization mismatch error between any two discretizations of the same manifold.
For completeness, we restate the full proposition:
Proposition A.1 (Full Statement).
Gauge-Invariant Spectral Self-Attention is a discretization-invariant Neural Operator with bounded discretization mismatch error. Let $M$ be a compact $m$-dimensional Riemannian manifold and $\mathcal{G}_{n}$ a sequence of graphs obtained by sampling $n$ nodes from $M$ with $n\to\infty$. Let $f$ be a spectral filter applied to a graph operator $P_{n}$ of $\mathcal{G}_{n}$, and let $\phi_{i}^{n}=f(P_{n})_{i\cdot}^{\top}$ be the spectral embedding of node $i$. Then:
  1. (i)
    (Special case: Laplacian eigenmaps.) When $f(\lambda)=\lambda^{-1/2}$ and $P_{n}=\mathcal{L}_{n}$, the inner products $\langle\phi_{i}^{n},\phi_{j}^{n}\rangle$ converge to the Green’s function $G_{M}(x_{i},x_{j})$ of the Laplace-Beltrami operator at rate $\mathcal{O}(n^{-1/(m+4)})$. More generally, for any Lipschitz spectral filter $f$, the kernel $K_{n}(i,j)=\langle\phi_{i}^{n},\phi_{j}^{n}\rangle=[f(P_{n})f(P_{n})^{\top}]_{ij}$ converges to a well-defined continuum kernel $K_{\infty}(x_{i},x_{j})$ at the same rate (Corollary A.3).
  2. (ii)
    (Discretization mismatch bound.) For any two discretizations $\mathcal{G}_{n}$ and $\mathcal{G}_{n^{\prime}}$ of the same manifold $M$ with $n\leq n^{\prime}$, the attention kernel values at corresponding points $x_{i},x_{j}\in M$ satisfy
    $\left|\langle\tilde{\phi}_{i}^{n},\tilde{\phi}_{j}^{n}\rangle-\langle\tilde{\phi}_{i}^{n^{\prime}},\tilde{\phi}_{j}^{n^{\prime}}\rangle\right|=\mathcal{O}\!\left(n^{-1/(m+4)}\right)+\mathcal{O}\!\left(r^{-1/2}\right),$
    where $\tilde{\phi}_{i}^{n}$ is the $i$-th row of $\widetilde{\Phi}^{n}=f(P_{n})\cdot R^{n}$. The two sources are independent: the first is the discretization mismatch error, the second the statistical estimation error from the random embedding.
  3. (iii)
    (Gauge invariance.) GIST’s attention kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ is exactly invariant under all gauge transformations of the eigenvector basis (sign flips, rotations within degenerate eigenspaces, solver choices). Combined with (ii), this ensures that learned parameters transfer across discretizations (discretization-invariance): the attention weights converge to a gauge-invariant continuum limit that does not depend on the arbitrary eigenvector conventions of any particular discretization.
The proof proceeds in three stages corresponding to the key technical components.
A.1.1 Stage 1: Spectral Convergence and Green’s Function
Proposition A.2.
Let $\mathcal{L}_{n}$ be the normalized graph Laplacian of a graph $\mathcal{G}_{n}$ obtained by sampling $n$ nodes from a compact $m$-dimensional Riemannian manifold $M$. Let $\lambda_{k}^{n},u_{k}^{n}$ be eigenvalues and eigenvectors of $\mathcal{L}_{n}$, and $\mu_{k},\psi_{k}$ the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta_{M}$ on $M$. Then:
  1. (a)
    Spectral convergence: $|\lambda_{k}^{n}-\mu_{k}|=\mathcal{O}(n^{-1/(m+4)})$ and $\|u_{k}^{n}-\psi_{k}\|_{L^{2}}=\mathcal{O}(n^{-1/(m+4)})$ (up to log factors).
  2. (b)
    Green’s function convergence: The inner products of Laplacian eigenmaps converge to the Green’s function:
    $\left\langle\phi_{i}^{n},\phi_{j}^{n}\right\rangle=\sum_{\lambda_{k}^{n}>0}\frac{1}{\lambda_{k}^{n}}(u_{k}^{n})_{i}(u_{k}^{n})_{j}\to G_{M}(x_{i},x_{j})+\mathcal{O}(n^{-1/(m+4)}),$
    where $G_{M}$ is the Green’s function of $\Delta_{M}$ on $M$.
Proof.
Part (a) follows from the spectral convergence theory for graph Laplacians on manifolds (CalderGarciaTrillos2020). With appropriate graph construction, both eigenvalues and eigenvectors of $\mathcal{L}_{n}$ converge to those of $\Delta_{M}$ at rate $n^{-1/(m+4)}$ (up to log factors). See CalderGarciaTrillos2020 for complete error estimates.
Part (b) follows from part (a) by the spectral theorem. The discrete Green’s function (pseudoinverse of the graph Laplacian) is given by the eigenfunction expansion:
$\langle\phi_{i}^{n},\phi_{j}^{n}\rangle$ $=\sum_{\lambda_{k}^{n}>0}\frac{1}{\lambda_{k}^{n}}(u_{k}^{n})_{i}(u_{k}^{n})_{j}.$
Using the convergence result from part (a), this sum converges to the continuum Green’s function $G_{M}(x_{i},x_{j})=\sum_{k=1}^{\infty}\frac{1}{\mu_{k}}\psi_{k}(x_{i})\psi_{k}(x_{j})$ at rate $\mathcal{O}(n^{-1/(m+4)})$. ∎
Corollary A.3 (Kernel convergence for general spectral filters).
Let $f$ be a spectral filter such that $g(\lambda):=f(\lambda)^{2}$ is Lipschitz on the spectrum of $\Delta_{M}$, with Lipschitz constant $L_{g}$. Then the discrete kernel $K_{n}(i,j)=\sum_{l}f(\lambda_{l}^{n})^{2}\,(u_{l}^{n})_{i}(u_{l}^{n})_{j}$ converges to the continuum kernel $K_{\infty}(x_{i},x_{j}):=\sum_{l}f(\mu_{l})^{2}\,\psi_{l}(x_{i})\psi_{l}(x_{j})$ at rate $\mathcal{O}(n^{-1/(m+4)})$.
Proof.
By the spectral convergence of the previous proposition (part a), $|\lambda_{k}^{n}-\mu_{k}|=\mathcal{O}(n^{-1/(m+4)})$ and $\|u_{k}^{n}-\psi_{k}\|_{L^{2}}=\mathcal{O}(n^{-1/(m+4)})$. The Lipschitz condition on $g$ gives $|g(\lambda_{k}^{n})-g(\mu_{k})|\leq L_{g}|\lambda_{k}^{n}-\mu_{k}|=\mathcal{O}(n^{-1/(m+4)})$. Combined with the eigenvector convergence (each term in the sum has both eigenvalue and eigenvector errors at the same rate), this yields $|K_{n}(i,j)-K_{\infty}(x_{i},x_{j})|=\mathcal{O}(n^{-1/(m+4)})$. When $f$ approximates $\lambda\mapsto 1/\sqrt{\lambda}$, $K_{\infty}$ reduces to the Green’s function $G_{M}$. ∎
A.1.2 Stage 2: Spectral Embeddings as Kernel Estimators
We consider the general class of approximate spectral embeddings $\widetilde{\Phi}=f(P)\cdot R\in\mathbb{R}^{N\times r}$, where $P$ is a graph operator, $f$ is a spectral filter (see §3 for definitions), and $R\in\mathbb{R}^{N\times r}$ is a random matrix with $\mathbb{E}[RR^{\top}]=\mathds{1}$. We write $\phi_{i}=f(P)_{i\cdot}^{\top}\in\mathbb{R}^{N}$ for the $i$-th row of $f(P)$ (the “exact” spectral embedding of node $i$) and $\tilde{\phi}_{i}\in\mathbb{R}^{r}$ for the $i$-th row of $\widetilde{\Phi}$ (its random projection); Laplacian eigenmaps equation 2 are the special case $f(\lambda)=\lambda^{-1/2}$, $P=\mathcal{L}$.
Lemma A.4 (Kernel estimator).
For any spectral filter $f$ and random matrix $R\in\mathbb{R}^{N\times r}$ with $\mathbb{E}[RR^{\top}]=\mathds{1}$, the inner products of $\widetilde{\Phi}=f(P)\cdot R$ satisfy
$\mathbb{E}\!\left[\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle\right]=K(i,j):=[f(P)f(P)^{\top}]_{ij}.$
When $P$ is symmetric with eigenvectors $u_{l}$, this reduces to $K(i,j)=\sum_{l}f(\lambda_{l})^{2}\,(u_{l})_{i}(u_{l})_{j}$.
Proof.
$\mathbb{E}[\widetilde{\Phi}\,\widetilde{\Phi}^{\top}]=f(P)\,\mathbb{E}[RR^{\top}]\,f(P)^{\top}=f(P)\cdot\mathds{1}\cdot f(P)^{\top}=f(P)f(P)^{\top}$. Taking the $(i,j)$ entry gives the result. ∎
Proposition A.5 (Exact gauge invariance of $K$).
The kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ is exactly invariant under all gauge transformations of the eigenvector basis: sign flips $u_{l}\to s_{l}u_{l}$, orthogonal rotations within degenerate eigenspaces, and different numerical solver choices.
Proof.
$K(i,j)=[f(P)f(P)^{\top}]_{ij}$ is an entry of the matrix $f(P)f(P)^{\top}$, which is a function of $P$ alone and does not depend on the choice of eigenvector basis. All gauge transformations leave $P$ unchanged, so $K$ is unchanged. ∎
Lemma A.6 (Concentration bound).
Let $\mathbf{a}=f(P)_{i\cdot}^{\top}$ and $\mathbf{b}=f(P)_{j\cdot}^{\top}$ be the $i$-th and $j$-th rows of $f(P)$, and let $R\in\mathbb{R}^{N\times r}$ have i.i.d. sub-Gaussian entries with mean zero and variance $1/r$. Then:
(i) Tail bound. By the Hanson-Wright inequality (Rudelson2013), for any $t>0$:
$\Pr\!\left[\left|\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle-K(i,j)\right|>t\right]\leq 2\exp\!\left(-c\cdot r\cdot\min\!\left(\frac{t^{2}}{\|\mathbf{a}\|^{2}\|\mathbf{b}\|^{2}},\frac{t}{\|\mathbf{a}\|\|\mathbf{b}\|}\right)\right),$
where $c>0$ is an absolute constant depending only on the sub-Gaussian norm of the entries of $R$.
(ii) Scaling of $r$. Fix a target accuracy $\varepsilon>0$ and a failure probability $\delta\in(0,1)$. Setting $t=\varepsilon\|\mathbf{a}\|\|\mathbf{b}\|$ in the bound above, and noting that for $\varepsilon\leq 1$ the minimum is achieved by the quadratic term $t^{2}/(\|\mathbf{a}\|^{2}\|\mathbf{b}\|^{2})=\varepsilon^{2}$, we require
$2\exp(-c\cdot r\cdot\varepsilon^{2})\leq\delta\qquad\Longleftrightarrow\qquad r\geq\frac{\log(2/\delta)}{c\,\varepsilon^{2}}.$
Therefore $r=\mathcal{O}(\varepsilon^{-2}\log(1/\delta))$ random-projection dimensions suffice to guarantee
$\left|\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle-K(i,j)\right|\leq\varepsilon\,\|\mathbf{a}\|\|\mathbf{b}\|$
with probability at least $1-\delta$. Equivalently, for fixed $\delta$, the estimation error is $\mathcal{O}(\|\mathbf{a}\|\|\mathbf{b}\|/\sqrt{r})$.
Proof.
Write $R=[\mathbf{r}_{1}\;\cdots\;\mathbf{r}_{r}]$ where each column $\mathbf{r}_{s}\in\mathbb{R}^{N}$ has i.i.d. sub-Gaussian entries with mean zero and variance $1/r$. Then
$\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle=\sum_{s=1}^{r}(\mathbf{a}^{\top}\mathbf{r}_{s})(\mathbf{b}^{\top}\mathbf{r}_{s})=\mathbf{a}^{\top}\!\left(\sum_{s=1}^{r}\mathbf{r}_{s}\mathbf{r}_{s}^{\top}\right)\!\mathbf{b}=\mathbf{a}^{\top}(RR^{\top})\mathbf{b}.$
Each summand $(\mathbf{a}^{\top}\mathbf{r}_{s})(\mathbf{b}^{\top}\mathbf{r}_{s})$ is independent with mean $\mathbb{E}[\mathbf{a}^{\top}\mathbf{r}_{s}\mathbf{r}_{s}^{\top}\mathbf{b}]=(1/r)\,\mathbf{a}^{\top}\mathbf{b}=(1/r)\,K(i,j)$, so $\mathbb{E}[\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle]=K(i,j)$. The product $(\mathbf{a}^{\top}\mathbf{r}_{s})(\mathbf{b}^{\top}\mathbf{r}_{s})$ can be written as $\mathbf{r}_{s}^{\top}(\mathbf{a}\mathbf{b}^{\top})\mathbf{r}_{s}$, a quadratic form in a sub-Gaussian random vector; summing $r$ such terms and applying the Hanson-Wright inequality (Rudelson2013) to $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle-K(i,j)=\mathbf{a}^{\top}(RR^{\top}-\mathds{1})\mathbf{b}$ yields the tail bound in part (i). Part (ii) follows by inverting this bound. ∎
A.1.3 Stage 3: Gauge Invariance of GIST
Proposition A.7.
GIST’s learned parameters do not depend on arbitrary gauge choices in the spectral decomposition, because GIST’s attention depends only on the exactly gauge-invariant kernel $K(i,j)$.
Proof.
GIST computes attention weights via linear attention with a feature map $\varphi$: the effective logits are $e_{ij}=\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle$. By Lemma A.4, $\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\rangle$ estimates $K(i,j)=[f(P)f(P)^{\top}]_{ij}$, and by Proposition A.5, $K(i,j)$ is exactly gauge-invariant. When $\varphi=\mathrm{ReLU}$ (as in Algorithm 2), the inner product $\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle$ converges to the arc-cosine kernel $k_{0}(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})$ as $r\to\infty$ (Cho2009), which depends only on the inner product $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}\to K(i,j)$ and is therefore gauge-invariant. All downstream computations operate on these gauge-invariant logits and on node features $x_{i}$ (which are gauge-independent), so the full architecture preserves gauge invariance. ∎
A.1.4 Discretization Mismatch Error Analysis
Combining all three stages, for two discretizations $\mathcal{G}_{n}$ and $\mathcal{G}_{n^{\prime}}$ of the same manifold $M$, the attention kernel mismatch is bounded by the triangle inequality:
[…]
where $K_{n}(i,j)=[f(P_{n})f(P_{n})^{\top}]_{ij}$ denotes the kernel on discretization $\mathcal{G}_{n}$. The middle term follows from Corollary A.3: both $K_{n}$ and $K_{n^{\prime}}$ converge to the same continuum kernel $K_{\infty}(x_{i},x_{j})=\sum_{l}f(\mu_{l})^{2}\psi_{l}(x_{i})\psi_{l}(x_{j})$ at rate $\mathcal{O}(n^{-1/(m+4)})$, so by the triangle inequality $|K_{n}(i,j)-K_{n^{\prime}}(i,j)|=\mathcal{O}(n^{-1/(m+4)})$. The total bound is therefore
$\left|\langle\tilde{\phi}_{i}^{n},\tilde{\phi}_{j}^{n}\rangle-\langle\tilde{\phi}_{i}^{n^{\prime}},\tilde{\phi}_{j}^{n^{\prime}}\rangle\right|=\mathcal{O}\!\left(\frac{1}{\sqrt{r}}\right)+\mathcal{O}\!\left(n^{-1/(m+4)}\right),$
with the two error sources being independent. Choosing $r=\mathcal{O}(\varepsilon^{-2})$ and $n=\mathcal{O}(\varepsilon^{-(m+4)})$ achieves total error $\varepsilon$.
Appendix B Algorithms
Note that in the pseudocode we use bold notation for matrices and vectors ($\mathbf{A},\mathbf{\Phi},\mathbf{Q}$) and follow the row-vector convention standard in machine learning: $\mathbf{\Phi}\in\mathbb{R}^{N\times r}$ has nodes as rows and embedding dimensions as columns. In the main text, we use non-bold notation for compactness, with $\phi_{i}\in\mathbb{R}^{r}$ representing individual column vectors and upper case characters denoting matrices.
Algorithm 1 Spectral Embeddings via FastRP (Chen2019e)
1:Graph adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, embedding dimensionality $r$, iteration power $k$
2:Matrix of $N$ node graph positional embeddings $\mathbf{\Phi}\in\mathbb{R}^{N\times r}$
3:Produce very sparse random projection $\mathbf{R}\in\mathbb{R}^{N\times r}$ according to Li2006
[…]
6:for $i=2$ to $k$ do
[…]
8:end for
[…]
10:return $\mathbf{\Phi}$
Below we provide pseudo-code for the core computations of GIST, the Gauge-Invariant Spectral Self-Attention block and the Gauge-Equivariant Spectral Self-Attention block. For illustration purposes, we compare the algorithms to a stripped down implementation of self-attention. Modifications relative to vanilla self-attention are highlighted in red (additions) and strikethrough (removals).
Algorithm 2 Gauge-Invariant Spectral Self-Attention
1:Node feature tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ , graph positional embeddings $\mathbf{\Phi}\in\mathbb{R}^{N\times r}$
2:Output sequence $\mathbf{O}\in\mathbb{R}^{N\times d}$ to be applied to features $\mathbf{X}$
3:// Compute attention matrices
[…]
7:// Compute linear attention with feature map $\varphi(x)=\mathrm{ReLU}(x)$
[…]
12:return $\mathbf{O}$
Algorithm 3 Gauge-Equivariant Spectral Self-Attention
1:Node feature tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ , graph positional embeddings $\mathbf{\Phi}\in\mathbb{R}^{N\times r}$
2:Output sequence $\mathbf{O}\in\mathbb{R}^{N\times{\color[rgb]{1,0,0}r}}$ to be applied to graph positional embeddings $\mathbf{\Phi}$
3:// Compute attention matrices
[…]
7:// Compute linear attention (Katharopoulos2020)
[…]
12:return $\mathbf{O}$
Note on the choice of Feature Map $\varphi$.
While in Algorithm 3 we do not need to impose that restriction, in Algorithm 2 we use the feature map $\varphi(x)=\text{ReLU}(x)$. This choice is theoretically motivated: when applied element-wise to random features, ReLU induces the arc-cosine kernel (Cho2009). Specifically, for vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{r}$, the inner product $\langle\varphi(\mathbf{a}),\varphi(\mathbf{b})\rangle$ converges (as $r\to\infty$) to a kernel function $k_{0}(\mathbf{a}^{\top}\mathbf{b})$ that depends only on the inner product $\mathbf{a}^{\top}\mathbf{b}$.
This property is crucial for preserving gauge invariance: the attention weights $\langle\varphi(\tilde{\phi}_{i}),\varphi(\tilde{\phi}_{j})\rangle\approx k_{0}(\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j})$ depend only on the inner product $\tilde{\phi}_{i}^{\top}\tilde{\phi}_{j}$, which by Lemma A.4 estimates the exactly gauge-invariant kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$. The resulting attention pattern therefore depends only on gauge-invariant quantities.
We note that the original linear attention work by Katharopoulos2020 used $\varphi(x)=\text{elu}(x)+1$. Empirically, this feature map also tends to work well in practice, and it is similar to ReLU in producing non-negative outputs. However, it is not known to correspond to any particular kernel function, and thus the theoretical guarantee of gauge invariance via kernel structure does not apply. Investigating other feature maps corresponding to different kernel functions (e.g., polynomial kernels, random Fourier features for RBF-like kernels) is left for future work.
Appendix C Additional Empirical Studies
C.1 GIST Hyperparameters robustness
In order to study the sensitivity of GIST’s performance to variations in its spectral embedding hyperparameters, we train multiple simplified GIST architectures (two-block Gauge-Invariant Spectral Self-Attention linear transformers) on the Cora benchmark while varying the power iteration parameter $k$ and the embedding dimension $r$ of the spectral embedding. These parameters directly control the quality of spectral embeddings while balancing computational efficiency.
As shown in Figure 2, GIST exhibits robust performance across a wide range of both parameters. The left panel sweeps $k$ with $r$ fixed, a clear but relatively shallow peak in accuracy around the optimal value of $k\approx 32$. This on one hand suggests that even modest iteration counts are sufficient to capture the essential spectral structure, but also indicates that the choice of the power iteration is quite robust. The right panel varies embedding dimension $r$ with $k$ fixed, demonstrating a smooth monotonic improvement as $r$ increases. Crucially, saturation occurs relatively quickly: performance gains beyond $r=256$ are marginal, validating our choice of reasonable embedding dimensions that maintain computational efficiency.
These results empirically validate two important properties: (1) GIST does not require extensive hyperparameter tuning around these spectral parameters, suggesting stable generalization; and (2) the linear end-to-end complexity achieved with modest $k$ and $r$ values is both computationally practical and empirically effective. Combined with the gauge-invariance guarantees that prevent dependence on arbitrary spectral choices, these hyperparameters provide a principled way and empirically robust way to control the approximation quality of spectral embeddings without sacrificing scalability.
Figure 2: Sensitivity study of GIST spectral embeddings parameters. The plots show the final test accuracy of a two-block Gauge-Invariant Spectral Self-Attention linear transformer trained on Cora while sweeping over the power iteration parameter $k$ with $r=256$ (left panel), and sweeping over the embedding dimension $r$ with $k=32$ (right panel). Test accuracy is fairly robust around the best value of either parameter. As expected, $r$ is monotonically related to higher performance, as higher $r$ correspond to better approximations of the eigenmaps. Accuracy conveniently saturate relatively fast, justifying the use of reasonably low $r$. The plots show mean test accuracy averaged across 10 seeds and corresponding standard deviation as error bars.
C.2 GIST Scalability Study
Figure 3: Scalability study of GIST. All experiments use a fixed 3-layer model while varying the hidden dimensionality. VRAM consumption was measured as a function of the number of nodes in the input graph. Graph sizes were controlled using random node dropout applied to samples from the Drivaernet dataset, enabling a systematic evaluation of memory scaling behavior.
From a computational standpoint, the end-to-end cost of GIST is dominated by two components: spectral embedding generation and the subsequent transformer blocks. For the former, we employ a FastRP-style approximation in which the Laplacian spectral information is captured via repeated multiplication of a sparse random walk matrix with a low-dimensional random projection. Each power iteration requires $\mathcal{O}(|E|r)$ operations, where $|E|$ is the number of edges and $r$ is the embedding dimension, and the total cost over $k$ iterations is $\mathcal{O}(k|E|r)$. On meshes and graphs with bounded average degree, $|E|=\mathcal{O}(N)$, so the overall spectral embedding stage scales linearly in the number of nodes $N$. This embedding is computed once per graph and then reused across all GIST layers, so its cost is amortized over the full network depth.
The GIST layers themselves preserve this linear scaling. Each block combines: (i) a feature branch based on linear attention, (ii) a local branch using graph convolution followed by linear attention, and (iii) a global branch using Gauge-Invariant and Gauge-Equivariant Spectral Self-Attention followed by linear attention. In all cases, attention is implemented in the form $\varphi(Q)(\varphi(K)^{\top}V)$ with an element-wise feature map $\varphi(\cdot)$, which yields $\mathcal{O}(Nd^{2})$ complexity for $d$-dimensional features instead of the $\mathcal{O}(N^{2}d)$ cost of quadratic attention. Together with the $\mathcal{O}(N)$ spectral embedding stage, this results in an overall complexity of $\mathcal{O}(N(d^{2}+rk))$ per forward pass, i.e., linear in the number of nodes. The empirical VRAM and wall-clock measurements in Figure 3 corroborate this analysis: both memory usage and forward time grow approximately linearly with the number of graph nodes for all hidden dimensions, up to graphs with hundreds of thousands of nodes sampled from DrivAerNet.
C.3 Multi-Scale Architecture Ablation Study
To validate the design choices in the Multi-Scale GIST architecture, we systematically ablate each of the three parallel branches shown in Figure 1 (right panel) on the PPI dataset, where the full architecture achieves state-of-the-art performance (see Table 3). For each ablation, we train with identical hyperparameters but remove one branch: (1) feature processing, (2) local graph convolution, or (3) global spectral attention. All experiments are repeated across 20 random seeds.
Table 6 shows that all three branches contribute meaningfully, with performance drops ranging from 4.29% to 7.90%. The local branch (Branch 2) has the strongest impact ($7.90\%$), validating the EfficientViT-inspired design principle that local operations provide focused information complementing the diffuse patterns from linear attention. The global spectral branch (Branch 3, $4.29\%$) confirms that long-range dependencies are essential, while the feature branch (Branch 1, $5.04\%$) provides complementary signal beyond structural information. Overall, these results demonstrate that the Multi-Scale GIST effectively integrates complementary information sources.
Table 6: Ablation study on PPI dataset showing the contribution of each branch of the Multi-Scale GIST (see Figure 1, right panel). Test accuracy is reported as percentage of baseline performance. Results averaged over 10 seeds with standard deviations indicated as uncertainty intervals.
Ablation Test Accuracy (% baseline) $\Delta$ (%)
No ablation 100.0 0.0
Branch 1 (feature) $95.0\pm 2.8$ 5.0-5.0
Branch 2 (local) $92.1\pm 1.8$ 7.9-7.9
Branch 3 (global) $95.7\pm 3.7$ 4.3-4.3
C.4 Pronunciation
GIST is pronounced with a soft G, like gist (the English word for the central point or essence of something), not with a hard G as in gift. Although insisting on a hard G won’t necessarily be considered wrong. After all, that is just a different gauge choice.